# PPO-ACT: Proximal Policy Optimization with Adversarial Curriculum Transfer for Spatial Public Goods Games

### v1

@@ Abstract
This study investigates cooperation evolution mechanisms in the spatial public goods game. A novel deep reinforcement learning framework, Proximal Policy Optimization with Adversarial Curriculum Transfer (PPO-ACT), is proposed to model agent strategy optimization in dynamic environments. Traditional evolutionary game models frequently exhibit limitations in modeling long-term decision-making processes. Deep reinforcement learning effectively addresses this limitation by bridging policy gradient methods with evolutionary game theory. Our study pioneers the application of proximal policy optimization’s continuous strategy optimization capability to public goods games through a two-stage adversarial curriculum transfer training paradigm. The experimental results show that PPO-ACT performs better in critical enhancement factor regimes. Compared to conventional standard proximal policy optimization methods, Q-learning and Fermi update rules, achieve earlier cooperation phase transitions and maintain stable cooperative equilibria. This framework exhibits better robustness when handling challenging scenarios like all-defector initial conditions. Systematic comparisons reveal the unique advantage of policy gradient methods in population-scale cooperation, i.e., achieving spatiotemporal payoff coordination through value function propagation. Our work provides a new computational framework for studying cooperation emergence in complex systems, algorithmically validating the punishment promotes cooperation hypothesis while offering methodological insights for multi-agent system strategy design.
@@
journal: arXiv
@@ 1 Introduction
Cooperation is a fundamental mechanism for sustaining and developing human societies and implies the foundational significance throughout civilizational evolution[1, 2, 3]. Collective behavior has driven human progress throughout history and continues to do so today. In early societies, cooperative efforts were central to hunting and agriculture. Modern civilization coordinates production and global cooperation. These patterns enhance environmental adaptation while accelerating knowledge growth and technological advancement. Contemporary global challenges such as climate change and public health crises further underscore the urgency of advancing cooperation mechanism research. Establishing sustainable cooperative paradigms amidst complex tensions between individual interests and collective welfare has become a central focus of interdisciplinary studies[4, 5]. Evolutionary game theory provides a systematic theoretical framework for studying the dynamic evolution of cooperative behavior[6, 7, 8, 9]. Among diverse applications, the public goods game (PGG) has become a cornerstone model for analyzing multi-agent cooperation dilemmas due to its precise characterization of tensions between individual and collective interests[10, 11, 12, 13]. This model reveals how cooperative behavior evolves in populations and provides key insights for solving real-world collective action problems.
@@
Traditional evolutionary game theory primarily relies on classical frameworks such as Fermi update rules[30] and replicator dynamics[31]. The classical frameworks effectively model immediate payoff effects and network influences in strategy diffusion. While demonstrating these capabilities, the approaches exhibit limitations in simulating key aspects of human decision-making complexity. A notable gap exists in modeling adaptive learning processes from historical experience and insufficient representation of strategic planning for long-term benefits. This gap has driven researchers toward the reinforcement learning framework[32, 33]. Reinforcement learning [34] creates a closed-loop learning system based on state-action-reward cycles. This allows agents to continuously improve their strategies in changing environments. It effectively handles long-term planning through value propagation while optimizing strategies with experience replay. Moreover, it dynamically maintains the balance between exploration and exploitation. This paradigm shift enables researchers to more accurately simulate real decision-makers trade-off processes between short-term benefits and long-term returns when analyzing cooperation strategy evolution[35, 36, 37].
@@
The Q-learning algorithm is widely used in evolutionary game theory because of its theoretical simplicity and practical effectiveness [38, 39]. It allows agents to adapt decisions based on historical experience and current environment states by building a Q-table to store state-action values[40, 41]. Notably, when applied to spatial PGG, Q-learning’s unique value iteration mechanism can maintain stable cooperative equilibrium even under free-riding incentives[42, 43, 44, 45]. Recent advances have further expanded Q-learning’s application dimensions. For instance, Yan et al.[46] innovatively combined periodic strategy updates with punishment mechanisms to establish an autonomous decision-making agent model. Similarly, Shen et al. [47] combined Q-learning with Fermi update rules. These advancements demonstrate how Q-learning provides new theoretical insights into the interaction between learning and imitation in spatial public goods games.
@@
While Q-learning’s tabular approach remains effective for small discrete action spaces, its reliance on value iteration becomes computationally prohibitive as strategy dimensions increase. In contrast, the Proximal Policy Optimization (PPO) algorithm [48] addresses these limitations through neural network-based policy parameterization, which not only circumvents Q-table memory constraints but also enables direct optimization of stochastic strategies through gradient ascent. Moreover, PPO’s clipped objective function inherently stabilizes policy updates, a critical advantage when coordinating multiple agents in high-dimensional discrete environments. Its reliable training stability and computational efficiency establish it as the benchmark for policy gradient methods. This dual-network architecture allows PPO to optimize policy parameters while ensuring training stability, leading to a more effective balance between exploration and exploitation. Unlike value-based methods that suffer from estimation bias in value functions, PPO directly optimizes policies to avoid suboptimal convergence. These characteristics make PPO particularly suitable for multi-agent scenarios requiring long-term policy planning. Recent research has significantly expanded PPO’s application domains. Sun et al.[49] innovatively integrated multi-attribute decision theory with PPO to address slow convergence in intelligent wargame training and agents’ low success rates with specific rules. Yu et al. [50] demonstrated PPO’s strong performance in cooperative multi-agent settings, achieving competitive or superior results in both final returns and sample efficiency. However, integrating modern reinforcement learning algorithms like PPO with evolutionary game theory still faces significant challenges. Current research has yet to fully uncover the diffusion dynamics of policy gradient methods in structured populations. The interaction effects between network topology and distributed learning processes remain insufficiently explored. These open questions provide promising directions for future research.
@@
We propose Proximal Policy Optimization with Adversarial Curriculum Transfer (PPO-ACT) for PGG. Our study pioneers the application of PPO in evolutionary game theory. The framework addresses cooperation evolution in spatial PGG through a novel two-stage training paradigm that builds upon curriculum learning [51] while introducing adversarial dynamics. It employs a dual-network architecture where the policy and value networks share underlying feature extraction layers. The framework’s two-stage curriculum learning design significantly enhances adaptability in complex environments. In Phase 1, cooperative foundations are established in high-reward conditions. In Phase 2, knowledge transfers to low-reward scenarios. This staged approach enables cooperators to better resist defectors under resource scarcity. Systematic simulation experiments validate the effectiveness of PPO-ACT. Results show this framework outperforms Q-learning and traditional evolutionary game methods by significantly enhancing cooperative behavior equilibrium levels and effectively suppressing free-riding diffusion. It shows strong adaptability to diverse initial conditions and changes in the environment. The Adversarial Curriculum Transfer (ACT) process facilitates quicker convergence by biasing the population toward cooperative strategies, outperforming random initialization approaches. Our research offers novel theoretical perspectives for understanding cooperation evolution mechanisms in social dilemmas through the lens of curriculum-based reinforcement learning. The methodological innovations also provide new analytical tools applicable to related studies in economics, ecology, and sociology.
@@ 2 Model
Consider a spatial PGG model defined on an $N=L\times{L}$ regular grid with periodic boundary conditions and von Neumann neighborhood ($k=4$). Each grid cell represents an agent, where agents form teams with their $k=4$ nearest neighbors for gameplay. Consider an agent set $\mathcal{A}=\{a_{1},...,a_{N}\}$ where each agent participates in $G=5$ overlapping game groups. Each group $\mathcal{G}_{i}$ is centered around an agent $i\in\mathcal{A}$, forming a local interaction neighborhood for evolutionary game dynamics. Define the strategy space $\mathcal{S}=\{C,D\}$, where $C$ denotes the cooperation strategy and $D$ denotes defection. The cooperation strategy contributes 1 unit to the public pool while the defection strategy contributes nothing. For any game group $g\in\mathcal{G}_{x}$, let $N_{C}^{g}$ denote the count of cooperators in the group, where each group consists of $k+1$ members. The resulting individual payoff function for the group is given by:
@@
where, $\Pi(s_{i}^{g})$ denotes the immediate payoff of agent $i$ in group $g$, $N_{C}^{g}=\sum_{j\in g}\mathbb{I}(s_{j}^{g}=C)$ is the number of cooperators in group $g$. $r>1$ is the enhancement factor. $s_{i}^{g}\in\{C,D\}$ denotes the strategy adopted by agent $i$ in group $g$. The cumulative payoff of agent $x$ is the sum of payoffs from all game groups it participates in:
@@ 2.1 PPO
where $\pi_{\theta}(a_{t}|s_{t})$ is the current policy representing the probability of selecting action $a_{t}$ at state $s_{t}$ in timestep $t$. $\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ denotes the old policy. $\theta$ represents the trainable parameter set of the Policy Network. In the PGG, the agent’s action $a_{t}$ is a binary contribution choice $c\in\{0,1\}$, where $c=1$ denotes cooperation and $c=0$ denotes defection. The state $s_{t}$ includes both the agent’s historical contribution record and the neighbors’ contribution information. The advantage function $A_{t}$ measures the relative benefit of taking action $a_{t}$ at state $s_{t}$, computed via generalized advantage estimation with the formula:
$A_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l},$ (5)
where $\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$ is the temporal difference error, representing the difference between the current reward at timestep $t$ and value function predictions. $r_{t}$ denotes the immediate reward obtained at timestep $t$. $\gamma\in[0,1)$ is the discount factor that determines the importance of future rewards. $\lambda\in[0,1)$ controls the weighting of future advantage estimations. $V(s_{t})$ is the state-value function representing the expected cumulative reward at state $s_{t}$, and $V(s_{t+1})$ is the state-value function for the next state $s_{t+1}$.
@@
$L^{PPO}(\theta)=L^{CLIP}(\theta)-\delta\cdot L^{VF}(\theta)+\rho\cdot L^{ENT}(\theta),$ (8)
@@
In PPO, the policy loss $\pi_{\theta}(a_{t}|s_{t})$ (Actor part) and value loss $V_{\theta}(s_{t})$ (Critic part) are computed by the network shown in Figure 1. The network takes the current state as input and shares a feedforward neural network layer for state feature extraction. This feedforward network consists of two fully connected layers with ReLU activation functions. The shared layer takes the input state $\mathbf{x}$ and outputs the hidden representation $\mathbf{h}$:
$\mathbf{h}=\text{ReLU}(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot x+\mathbf{b}_{1})+\mathbf{b}_{2}),$ (9)
@@
Figure 1: The Actor-Critic network architecture in the PPO algorithm
@@
$\mathbf{a}=\mathbf{W}_{\text{actor}}\cdot\mathbf{h}+\mathbf{b}_{\text{actor}},$ (10)
@@
$\pi(a|s)=\text{softmax}(\mathbf{a}).$ (11)
@@
$V(s)=\mathbf{W}_{critic}\cdot\mathbf{h}+\mathbf{b}_{critic},$ (12)
@@ 2.2 PPO-ACT
$R_{t}=\Pi_{x},$ (13)
where $\Pi_{x}$ is the cumulative payoff of agent $x$, calculated by:
$\Pi_{x}=\sum^{G}_{g=1}\Pi_{x}^{g},$ (14)
with $\Pi_{x}^{g}$ being agent $x$’s payoff in group $g$.
@@
5:
6:Phase 1: Cooperative Policy Initialization
7:Initial enhancement factor $r_{1}$
8:for $t=1$ to $T_{1}$ do
9:     for each agent $i$ do
10:         Select action $a_{t}$ according to current policy $\pi_{\theta}$
11:         Execute PGG, compute payoff $\Pi_{i}$ and reward $\mathcal{R}_{t}$
12:     end for
13:     Standard PPO updates:
14:      Compute advantage function $A_{t}$ and target value $V_{t}^{\text{target}}$
15:      Compute PPO objective function $L^{PPO}(\theta)$
16:      Update policy network $\pi_{\theta}$ and value network $V_{\theta}$ via gradient descent
17:end for
18:
19:Phase 2: Adversarial Curriculum Transfer
20:Initial enhancement factor $r_{2}$
21:for $t=1$ to $T_{2}$ do
22:     for each agent $i$ do
23:         Select action $a_{t}$ according to current policy $\pi_{\theta}$
24:         Execute PGG, compute payoff $\Pi_{i}$ and reward $\mathcal{R}_{t}$
25:     end for
26:     Standard PPO updates:
27:      Compute advantage function $A_{t}$ and target value $V_{t}^{\text{target}}$
28:      Compute PPO objective function $L^{PPO}(\theta)$
29:      Update policy network $\pi_{\theta}$ and value network $V_{\theta}$ via gradient descent
30:end for

### v2

@@ Abstract
This study investigates cooperation evolution mechanisms in the spatial public goods game. A novel deep reinforcement learning framework, Proximal Policy Optimization with Adversarial Curriculum Transfer (PPO-ACT), is proposed to model agent strategy optimization in dynamic environments. Traditional evolutionary game models often exhibit limitations in modeling long-term decision-making processes. Imitation-based rules (e.g., Fermi) lack strategic foresight, while tabular methods (e.g., Q-learning) fail to capture spatial-temporal correlations. Deep reinforcement learning effectively addresses these limitation by bridging policy gradient methods with evolutionary game theory. Our study pioneers the application of proximal policy optimization’s continuous strategy optimization capability to public goods games through a two-stage adversarial curriculum transfer training paradigm. The experimental results show that PPO-ACT performs better in critical enhancement factor regimes. Compared to conventional standard proximal policy optimization methods, Q-learning and Fermi update rules, achieve earlier cooperation phase transitions and maintain stable cooperative equilibria. This framework exhibits better robustness when handling challenging scenarios like all-defector initial conditions. Systematic comparisons reveal the unique advantage of policy gradient methods in population-scale cooperation, i.e., achieving spatiotemporal payoff coordination through value function propagation. Our work provides a new computational framework for studying cooperation emergence in complex systems, algorithmically validating the punishment promotes cooperation hypothesis while offering methodological insights for multi-agent system strategy design.
@@
journal: Chaos, Solitons & Fractals11footnotetext: https://github.com/geek12138/PPO-ACT
@@
[…]
[5]organization=State Key Laboratory of Public Big Data, College of Big Data and Information Engineering, addressline=Guizhou University, city=Guiyang, postcode=550025, state=Guizhou, country=China
@@ 1 Introduction
Cooperation is a fundamental mechanism for sustaining and developing human societies and implies the foundational significance throughout civilizational evolution [1, 2, 3]. Collective behavior has driven human progress throughout history and continues to do so today. In early societies, cooperative efforts were central to hunting and agriculture. Modern civilization coordinates production and global cooperation. These patterns enhance environmental adaptation while accelerating knowledge growth and technological advancement. Contemporary global challenges such as climate change and public health crises further underscore the urgency of advancing cooperation mechanism research. Establishing sustainable cooperative paradigms amidst complex tensions between individual interests and collective welfare has become a central focus of interdisciplinary studies [4, 5]. Evolutionary game theory provides a systematic theoretical framework for studying the dynamic evolution of cooperative behavior [6, 7, 8]. Among diverse applications, the public goods game (PGG) has become a cornerstone model for analyzing multi-agent cooperation dilemmas due to its precise characterization of tensions between individual and collective interests [9, 10, 11]. It is worth noting that Duh et al. [12] investigated public goods games on random geometric graphs in hyperbolic spaces. This model reveals how cooperative behavior evolves in populations and provides key insights for solving real-world collective action problems.
@@
Traditional evolutionary game theory primarily relies on classical frameworks such as Fermi update rules [29] and replicator dynamics [30]. The classical frameworks effectively model immediate payoff effects and network influences in strategy diffusion. While demonstrating these capabilities, the approaches exhibit limitations in simulating key aspects of human decision-making complexity. A notable gap exists in modeling adaptive learning processes from historical experience and insufficient representation of strategic planning for long-term benefits. This gap has driven researchers toward the reinforcement learning framework [31, 32]. Reinforcement learning [33] creates a closed-loop learning system based on state-action-reward cycles. This allows agents to continuously improve their strategies in changing environments. It effectively handles long-term planning through value propagation while optimizing strategies with experience replay. Moreover, it dynamically maintains the balance between exploration and exploitation. Recent advances in reinforcement learning have enabled more sophisticated modeling of emergent cooperation. Phan et al. [34] proposed a mutual acknowledgment token exchange that implements emergent cooperation via a reinforcement learning-based token exchange, achieving high fault tolerance in multi-agent systems. This paradigm shift enables researchers to more accurately simulate real decision-makers trade-off processes between short-term benefits and long-term returns when analyzing cooperation strategy evolution [35, 36, 37].
@@
The Q-learning algorithm is widely used in evolutionary game theory because of its theoretical simplicity and practical effectiveness [38, 39]. It allows agents to adapt decisions based on historical experience and current environment states by building a Q-table to store state-action values [40, 41]. Notably, when applied to spatial PGG, Q-learning’s unique value iteration mechanism can maintain stable cooperative equilibrium even under free-riding incentives [42, 43, 44]. Recent advances have further expanded Q-learning’s application dimensions. For instance, Yan et al. [45] innovatively combined periodic strategy updates with punishment mechanisms to establish an autonomous decision-making agent model. Similarly, Shen et al. [46] combined Q-learning with Fermi update rules. These advancements demonstrate how Q-learning provides new theoretical insights into the interaction between learning and imitation in spatial public goods games.
@@
While Q-learning’s tabular approach remains effective for small discrete action spaces, its reliance on value iteration becomes computationally prohibitive as strategy dimensions increase. In contrast, the Proximal Policy Optimization (PPO) algorithm [47] addresses these limitations through neural network-based policy parameterization, which not only circumvents Q-table memory constraints but also enables direct optimization of stochastic strategies through gradient ascent. Moreover, PPO’s clipped objective function inherently stabilizes policy updates, a critical advantage when coordinating multiple agents in high-dimensional discrete environments. Its reliable training stability and computational efficiency establish it as the benchmark for policy gradient methods. This dual-network architecture allows PPO to optimize policy parameters while ensuring training stability, leading to a more effective balance between exploration and exploitation. Unlike value-based methods that suffer from estimation bias in value functions, PPO directly optimizes policies to avoid suboptimal convergence. These characteristics make PPO particularly suitable for multi-agent scenarios requiring long-term policy planning. Contemporary advances [48, 49, 50] in multi-agent reinforcement learning have yielded substantial theoretical and empirical evidence supporting its efficacy in complex cooperative task domains. Building on these foundations, Sun et al. [51] innovatively integrated multi-attribute decision theory with PPO to address slow convergence in intelligent wargame training and agents’ low success rates with specific rules. Yu et al. [52] demonstrated PPO’s strong performance in cooperative multi-agent settings, achieving competitive or superior results in both final returns and sample efficiency. However, integrating modern reinforcement learning algorithms like PPO with evolutionary game theory still faces significant challenges. Current research has yet to fully uncover the diffusion dynamics of policy gradient methods in structured populations. The interaction effects between network topology and distributed learning processes remain insufficiently explored. These open questions provide promising directions for future research.
@@
We propose Proximal Policy Optimization with Adversarial Curriculum Transfer (PPO-ACT) for PGG. Our study pioneers the application of PPO in evolutionary game theory. The framework addresses cooperation evolution in spatial PGG through a novel two-stage training paradigm that builds upon curriculum learning [53] while introducing adversarial dynamics. Curriculum learning is used to enhance tasks based on reinforcement learning, such as robot navigation [54], unmanned aerial vehicles collaborative work [55]. It employs a dual-network architecture where the policy and value networks share underlying feature extraction layers. The framework’s two-stage curriculum learning design significantly enhances adaptability in complex environments. In Phase 1, cooperative foundations are established in high-reward conditions. In Phase 2, knowledge transfers to low-reward scenarios. This staged approach enables cooperators to better resist defectors under resource scarcity. Systematic simulation experiments validate the effectiveness of PPO-ACT. Results show this framework outperforms Q-learning and traditional evolutionary game methods by significantly enhancing cooperative behavior equilibrium levels and effectively suppressing free-riding diffusion. It shows strong adaptability to diverse initial conditions and changes in the environment. The Adversarial Curriculum Transfer (ACT) process facilitates quicker convergence by biasing the population toward cooperative strategies, outperforming random initialization approaches. Our research offers novel theoretical perspectives for understanding cooperation evolution mechanisms in social dilemmas through the lens of curriculum-based reinforcement learning. The methodological innovations also provide new analytical tools applicable to related studies in economics, ecology, and sociology.
@@ 2 Model
Consider a spatial PGG model defined on an $N=L\times{L}$ regular grid with periodic boundary conditions and von Neumann neighborhood ($k=4$). Each grid cell represents an agent, where agents form teams with their $k=4$ nearest neighbors for gameplay. Consider an agent set $\mathcal{\hat{A}}=\{\hat{a}_{1},...,\hat{a}_{N}\}$ where each agent participates in $G=5$ overlapping game groups. Each group $\mathcal{G}_{i}$ is centered around an agent $i\in\mathcal{\hat{A}}$, forming a local interaction neighborhood for evolutionary game dynamics. Define the strategy space $\mathcal{S}=\{C,D\}$, where $C$ denotes the cooperation strategy and $D$ denotes defection. The cooperation strategy contributes 1 unit to the public pool while the defection strategy contributes nothing. For any game group $g\in\mathcal{G}_{x}$, let $N_{C}^{g}$ denote the count of cooperators in the group, where each group consists of $k+1$ members. The resulting individual payoff function for the group is given by:
@@
where, $\Pi(s_{i}^{g})$ denotes the immediate payoff of agent $i$ in group $g$, $N_{C}^{g}=\sum_{j\in g}\mathbb{I}(s_{j}^{g}=C)$ is the number of cooperators in group $g$. $r>1$ is the enhancement factor. $s_{i}^{g}\in\{C,D\}$ denotes the strategy adopted by agent $i$ in group $g$. The cumulative payoff of agent $i$ is the sum of payoffs from all game groups it participates in:
@@ 2.1 PPO
where $\pi_{\theta}(a_{t}|s_{t})$ is the current policy representing the probability of selecting action $a_{t}$ at state $s_{t}$ in timestep $t$. $\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ denotes the old policy. $\theta$ represents the trainable parameter set of the Policy Network. In the PGG, the agent’s action $a_{t}$ is a binary contribution choice $\hat{c}\in\{0,1\}$, where $\hat{c}=1$ denotes cooperation and $\hat{c}=0$ denotes defection. The state $s_{t}$ includes both the agent’s historical contribution record and the neighbors’ contribution information. The advantage function $A_{t}$ measures the relative benefit of taking action $a_{t}$ at state $s_{t}$, computed via generalized advantage estimation with the formula:
$A_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\psi_{t+l},$ (5)
where $\psi_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$ is the temporal difference error, representing the difference between the current reward at timestep $t$ and value function predictions. $r_{t}$ denotes the immediate reward obtained at timestep $t$. $\gamma\in[0,1)$ is the discount factor that determines the importance of future rewards. $\lambda\in[0,1)$ controls the weighting of future advantage estimations. $V(s_{t})$ is the state-value function representing the expected cumulative reward at state $s_{t}$, and $V(s_{t+1})$ is the state-value function for the next state $s_{t+1}$.
@@
$L^{PPO}(\theta)=L^{CLIP}(\theta)-\delta\cdot L^{VF}(\theta)+\rho\cdot L^{ENT}(\theta),\vspace{-1mm}$ (8)
@@
In PPO, the policy loss $\pi_{\theta}(a_{t}|s_{t})$ (Actor part) and value loss $V_{\theta}(s_{t})$ (Critic part) are computed by the network shown in Figure 1. The network takes the current state as input and shares a feedforward neural network layer for state feature extraction. This feedforward network consists of two fully connected layers with ReLU [56] activation functions. The shared layer takes the input state $\mathbf{x}$ and outputs the hidden representation $\mathbf{h}$:
$\mathbf{h}=\text{ReLU}(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot x+\mathbf{b}_{1})+\mathbf{b}_{2}),\vspace{-1mm}$ (9)
@@
Figure 1: The PPO-ACT framework employs an Actor-Critic architecture where a shared feature extractor processes environmental states, enabling simultaneous optimization of both policy distribution $\pi(a|s)$ through the Actor pathway and state-value estimation $V(s)$ via the Critic pathway, with parameter sharing between these components significantly enhancing spatial correlation learning in PGG.
@@
$\mathbf{a}=\mathbf{W}_{\text{actor}}\cdot\mathbf{h}+\mathbf{b}_{\text{actor}},\vspace{-1mm}$ (10)
@@
$\pi(a|s)=\text{softmax}(\mathbf{a}).\vspace{-1mm}$ (11)
@@
$V(s)=\mathbf{W}_{critic}\cdot\mathbf{h}+\mathbf{b}_{critic},\vspace{-1mm}$ (12)
@@ 2.2 PPO-ACT
$R_{t}=\Pi_{i},$ (13)
where $\Pi_{i}$ is the cumulative payoff of agent $i$, calculated by:
$\Pi_{i}=\sum^{G}_{g=1}\Pi_{i}^{g},$ (14)
with $\Pi_{i}^{g}$ being agent $i$’s payoff in group $g$.
@@
5:Phase 1: Cooperative Policy Initialization
6:Initial enhancement factor $r_{1}$
7:for $t=1$ to $T_{1}$ do
8:     for each agent $i$ do
9:         Select action $a_{t}$ according to current policy $\pi_{\theta}$
10:         Execute PGG, compute payoff $\Pi_{i}$ and reward $\mathcal{R}_{t}$
11:     end for
12:     Standard PPO updates:
13:      Compute advantage function $A_{t}$ and target value $V_{t}^{\text{target}}$
14:      Compute PPO objective function $L^{PPO}(\theta)$
15:      Update policy network $\pi_{\theta}$ and value network $V_{\theta}$ via gradient descent
16:end for
17:Phase 2: Adversarial Curriculum Transfer
18:Initial enhancement factor $r_{2}$
19:for $t=1$ to $T_{2}$ do
20:     for each agent $i$ do
21:         Select action $a_{t}$ according to current policy $\pi_{\theta}$
22:         Execute PGG, compute payoff $\Pi_{i}$ and reward $\mathcal{R}_{t}$
23:     end for
24:     Standard PPO updates:
25:      Compute advantage function $A_{t}$ and target value $V_{t}^{\text{target}}$
26:      Compute PPO objective function $L^{PPO}(\theta)$
27:      Update policy network $\pi_{\theta}$ and value network $V_{\theta}$ via gradient descent
28:end for